# Learning when to rank: Estimation of partial rankings from sparse, noisy comparisons

### v2

@@ Appendix C Tau Recovery
Figure 7: Correlations of inferred rankings with planted rankings. (a) $\tau_{B}$ scores between the rankings inferred by the BT and PR models with respect to the ground truth ranking for the case in which there are no partial rankings ($R=N=150$). (b) $\tau_{B}$ scores between the rankings inferred by the BT and PR models with respect to the ground truth ranking for the case in which there are three planted partial rankings ($R=3$). All results are averaged over 10 different simulations and error bars represent 2 standard errors in the mean.
@@
In Fig. 7, we show Kendall’s $\tau_{B}$ rank correlation coefficient as a function of the average degree for synthetically generated networks consisting of $N=150$ nodes. Panel (a) corresponds to the case $R=N=150$, where each node is assigned a unique ranking, implying no partial rankings are present. Panel (b) illustrates the scenario where $R=3$ unique partial rankings are imposed, with the nodes distributed at random among these three ranks.
@@
In the case where no partial rankings are present (panel (a)), we observe that the BT model consistently outperforms the PR model in terms of rank correlation with the ground truth rankings, with the PR model obtaining similar $\langle\tau_{B}\rangle$ scores only at very high values of $\langle k\rangle$. Note that, as shown in Fig. 2(d), the PR model never recovers the correct number of rankings, even in dense regimes. However, the PR model is able to increase the inferred number of partial rankings as more data becomes available, which improves its rank correlation with the ground truth. By the time the PR model infers approximately $R=50$ rankings, it achieves $\tau_{B}$ scores of roughly 0.99.
@@
When partial rankings are present (panel (b)), the $\langle\tau_{B}\rangle$ value for the BT model plateaus around 0.8. This behaviour likely stems from the BT model’s inherent assumption of a complete ranking, which limits its ability to adapt once it has achieved the best possible alignment it can with the ground truth rankings. On the other hand, the PR model is able to obtain perfect correlation already at relatively low values of $\langle k\rangle$, thanks to its ability to account for and adapt to the presence of partial rankings in the data.
Appendix D Comparison with 1D Clustering
@@
Figure 8: Comparison of PR with 1D clustering of BT scores. (a) Number of ranks inferred by the partial rankings algorithm and mean shift clustering as a function of the average degree of the network for the case where there are no partial rankings ($R=N=150$). (b) Number of ranks inferred by the partial rankings algorithm and mean shift clustering as a function of the average degree of the network for the case in which there are three planted rankings ($R=3$, indicated by the solid black line). Results are averaged over $10$ different network realisations for each $\langle k\rangle$ value and error bars indicate 2 standard errors in the mean.

### v3

@@ Appendix D Tau Recovery
Figure 10: Correlations of inferred rankings with planted rankings. (a) $\tau_{B}$ scores between the rankings inferred by the BT and PR models with respect to the ground truth ranking for the case in which there are no partial rankings ($R=N=150$). (b) $\tau_{B}$ scores between the rankings inferred by the BT and PR models with respect to the ground truth ranking for the case in which there are three planted partial rankings ($R=3$). All results are averaged over 10 different simulations and error bars represent 2 standard errors from the mean.
Figure 11: Inference of planted scores. (a) Histograms of planted ground truth scores, BT inferred scores, and PR inferred scores. (b) BT and PR inferred scores, ordered by increasing player strength. The horizontal black lines correspond to the ground-truth planted scores. All results were obtained for $N=50$ nodes and $R=5$ planted partial rankings
@@
In Fig. 10, we show Kendall’s $\tau_{B}$ rank correlation coefficient as a function of $\langle m\rangle$ for synthetically generated networks consisting of $N=150$ nodes. Panel (a) corresponds to the case $R=N=150$, where each node is assigned a unique ranking, implying no partial rankings are present. Panel (b) illustrates the scenario where $R=3$ unique partial rankings are imposed, with the nodes distributed at random among these three ranks.
@@
In the case where no partial rankings are present (panel (a)), we observe that the BT model consistently outperforms the PR model in terms of rank correlation with the ground truth rankings, with the PR model obtaining similar $\langle\tau_{B}\rangle$ scores only at very high values of $\langle m\rangle$. Note that, as shown in Fig. 3, the PR model never recovers the correct number of rankings, even in dense regimes. However, the PR model is able to increase the inferred number of partial rankings as more data becomes available, which improves its rank correlation with the ground truth. By the time the PR model infers approximately $R=50$ rankings, it achieves $\tau_{B}$ scores of roughly 0.99.
@@
When partial rankings are present (panel (b)), the $\langle\tau_{B}\rangle$ value for the BT model plateaus around 0.8. This behaviour likely stems from the BT model’s inherent assumption of a complete ranking, which limits its ability to adapt once it has achieved the best possible alignment it can with the ground truth rankings. On the other hand, the PR model is able to obtain perfect correlation already at relatively low values of $\langle m\rangle$, thanks to its ability to account for and adapt to the presence of partial rankings in the data.
Fig. 11 shows a comparison of the player scores inferred by both the BT and PR models with respect to the ground-truth planted scores for the case $N=50$, $R=5$. We observe that both the BT and PR models are able to recover scores which are very close to the planted ones. However, the BT model is unable to appropriately account for small fluctuations in match outcomes and assigns each node a unique score, even when these are very close. On the other hand, the PR algorithm is able to leverage the coarse-grained structure of the scores to group players into the correct number of partial rankings.
Appendix E Comparison with 1D Clustering
@@
Figure 12: Comparison of PR with 1D clustering of BT scores. (a) Number of ranks inferred by the partial rankings algorithm and mean shift clustering as a function of $\langle m\rangle$ for the case where there are no partial rankings ($R=N=150$). (b) Number of ranks inferred by the partial rankings algorithm and mean shift clustering as a function of $\langle m\rangle$ for the case in which there are three planted rankings ($R=3$, indicated by the solid black line). Results are averaged over 1010 different network realizations for each $\langle m\rangle$ value, and error bars indicate two standard errors from the mean.